[Ncep.nhc.nco_contacts] TSB Morning Rounds - Sat Sep 09, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Sat Sep 9 14:33:11 UTC 2017


Here is today's summary of NHC computer operations:
--- Mello

----------

TAFB:
1. NFC has an open ticket to monitor GFE in storm mode.
HSU:
1. Eric Blake will issue some products from WPC today.
2. Melllo will work a radar shift this afternoon.
3. Major comms issue with Katia Recon mission. The problem was external to  
NHC and appears to be an isolated incident.
Other items:
1. Lockdown will occur at 7:00 PM today.
2. Rain bands caused many trees limbs to fall on roads in Kendall.
NHC RFCs:
No RFC during the current storm

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170909/698cf78b/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list